OpenStudy (anonymous):

simplify 6√28 - 3√63 + 9√7

OpenStudy (anonymous):

\[6*2\sqrt{7} - 3*3\sqrt{7} + 9\sqrt{7} = (12 - 9 + 9)\sqrt{7} = 12\sqrt{7}\]